find the prime
27,29,33,35

Answer by jim_thompson5910(35256)

27 = 9*3 .... so 27 is NOT prime

33 = 3*11 ... so 33 is NOT prime

35 = 7*5 .... so 35 is NOT prime

By process of elimination and by the fact that 29 is indeed prime (the only factors are 29 and 1), the answer is 29.